What is the spring constant k if a spring stretches .5m when a 10 kg mass is suspended from it?
F = k x
1
Expert's answer
2020-12-10T11:02:06-0500

We can find the spring constant from the Hooke's Law:


"F=kx,""mg=kx,""k=\dfrac{mg}{x}=\dfrac{10\ kg\cdot 9.8\dfrac{m}{s^2}}{0.5\ m}=196\ \dfrac{N}{m}."

Answer:

"k=196\ \dfrac{N}{m}."